# Determinantal random subgraphs

### v3

@@ 15.4. Scaling limits
We could also consider the case of a discretization of a curved disk by a fine graph: for example, take a portion of the round sphere, discretize the $1$-form $dF$, where $F$ is the curvature $2$-form and consider the determinantal random subgraph in $\mathcal{C}_{1}(\mathsf{G})$ defined by $\mathop{\mathrm{im}}d\oplus\mathbb{K}F$. Does this random spanning subgraph, or the unique random simple loop it defines, have a scaling limit?
Before tackling two-dimensional scaling limits, one could also consider simpler one-dimensional limits. For instance, we could consider the Laplacian spanning forest on a cycle graph with $n$ edges. Does the process associated with the kernel $\mathsf{k}$, or more likely, the process associated with the kernel $\mathsf{k}^{a}$ for $10<a<$ 1 (where $a$ is chosen appropriately with $n$), have a scaling limit on the unit circle when $n$ tends to infinity?

### v4

@@ 15.4. Scaling limits
We could also consider the case of a discretization of a curved disk by a fine graph: for example, take a portion of the round sphere, discretize the $1$-form dFdF, where $F$ is the curvature $2$-form and consider the determinantal random subgraph in $\mathcal{C}_{1}(\mathsf{G})$ defined by $\mathop{\mathrm{im}}d\oplus\mathbb{K}F$. Does this random spanning subgraph, or the unique random simple loop it defines, have a scaling limit?
Before tackling two-dimensional scaling limits, one could also consider simpler one-dimensional limits. For instance, we could consider the Laplacian spanning forest on a cycle graph with $n$ edges. Does the process associated with the kernel $\mathsf{k}$, or more likely, the process associated with the kernel $\mathsf{k}^{a}$ for $0<a<1$ (where $a$ is chosen appropriately with $n$), have a scaling limit on the unit circle when $n$ tends to infinity?